please help me solve this problem
The hypotenuse side c of a triangle is 13 inches long. Which of the following pairs of measurements could be correct for the lengths of the other two sides of the triangle. Note A2 + B2 = C2

Answer by Alan3354(69443)

please help me solve this problem
The hypotenuse side c of a triangle is 13 inches long. Which of the following pairs of measurements could be correct for the lengths of the other two sides of the triangle. Note A2 + B2 = C2
--------------------
Nothing followed.
Use lower case for sides, Upper Case for angles.
Use ^ (Shift 6) for exponents.
a^2 + b^2 = c^2
-----
PS Do you think we don't know that?